# Dynamical friction and measurements of the splashback radius in galaxy clusters

### v1

@@ 1 Introduction
In this work we use data from the IllustrisTNG (TNG) cosmological simulations to reconstruct simplified realizations of clusters and subhaloes before evolving the subhalo orbits further forwards in time. We use cluster information from TNG because, although we create an idealized simulation to test DF, the initial phase-space distribution of the subhaloes is crucial for replicating the splashback radius. Creating a realistic phase-space distribution of subhaloes organically is very difficult, and it is easy to inadvertantly destroy the splashback radius, which reflects billions of years of the cluster’s evolution. These complications are why we use the TNG data instead of creating our own clusters, even if that approach could, in theory, be more flexible. The next sections introduce TNG (2.1) and how (sub)haloes in TNG are identified and selected (2.2), before describing the details of our simulation (2.3) and fitting procedure (2.4).
@@ 2.1 IllustrisTNG Simulations
We utilize the $z=0$ snapshot of the largest-volume IllustrisTNG cosmological simulations, as introduced by Nelson et al. (2018); Pillepich et al. (2018a) and publicly released in Nelson et al. (2019). Though the full TNG simulation series includes three box sizes (approximately 50, 100, and 300 Mpc side lengths), we use the higest-resolution of the largest volume runs, labeled as TNG300-1. TNG300-1 has a periodic box with side length 302 Mpc, providing a large sample of galaxy clusters. This volume contains $2500^{3}$ dark matter particles with masses of $5.9\times 10^{7}~{}\mathrm{M_{\odot}}$, and $2500^{3}$ gas particles with target cell masses of $1.1\times 10^{7}~{}\mathrm{M_{\odot}}$. The dark matter particles have a gravitational softening length of 1.5 kpc in comoving units for $z>1$ and physical units for $z\leq 1$, whereas the gas cells utilize an adaptive softening length (minimum 0.37 kpc).
TNG is built on the magnetohyrodynamical (MHD) moving-mesh code Arepo (Springel, 2010), with an updated version of the Illustris galaxy formation model (Vogelsberger et al., 2013). In the model, gas cells cool through radiative processes and metal line cooling, until they stochastically form stars. As the stars evolve, stellar winds and supernovae return mass, energy, and enriched metals to the gas. Black holes are also included in the simulation; they are seeded in haloes with masses of $1.2\times 10^{6}~{}\rm M_{\odot}$. TNG updates parts of the Illustris model, including radio mode feedback from active galactic nuclei (AGN) (Weinberger et al., 2017), new supernova wind model (Pillepich et al., 2018a), and also includes improvements to the numerical convergence of Arepo (Pakmor et al., 2016). See Weinberger et al. (2017) for more information about other AGN feedback modes.
IllustrisTNG follows a cosmology consistent with the Planck Collaboration et al. (2016) results. All the runs use cosmological parameters of $\Omega_{m}=\Omega_{dm}+\Omega_{b}=0.3089,\Omega_{b}=0.0486,~{}\rm and~{}\Omega_{\Lambda}=0.6911$. Additionally, we use $\sigma_{8}=0.8159,n_{s}=0.9667$. The Hubble constant is set as $H_{0}=100h$ km s-1 Mpc-1, with $h=0.6774$ (Planck Collaboration et al., 2016). Some runs are dark-matter (gravity) only, and the rest are fully hydrodynamical runs. We utilize the hydrodynamical runs.
@@ 2.2 Halo selection and galaxy definition
We follow the halo selection criteria of O’Neil et al. (2021), choosing all haloes of $M_{\mathrm{200,mean}}>10^{13}\mathrm{M_{\odot}}$. Beyond the mass requirement, O’Neil et al. (2022) additionally remove haloes that are within $10R_{\mathrm{200,mean}}$ of a more massive halo, preferentially selecting haloes that are isolated and have not been disrupted. This filter is important given that we care about the density profiles outside of the cluster center, and because further simulations of the cluster (see Section 2.3) are based on isolated spherical potentials for each halo. From this process we were left with 1401 haloes at $z=0$. However, we further filter the most massive halo because we identify its density profile as having been significantly disrupted by some less massive haloes within $10R_{\mathrm{200,mean}}$, producing a final sample of 1400 haloes.
@@ 2.3 Simulation details
The subhaloes are integrated in the static gravitational potential of the cluster for 10 Gyr, with the subhalo positions outputted every 1 Gyr. The simulation was run twice from the same starting conditions. In one case, we run with the gravity from a static cluster potential as the only force, in which case there is no dynamical friction implemented. In the second, we run the galpy integration with both the cluster potential and a dynamical friction force, which is described in the next section.
@@ 2.3.1 Analytical implementation of dynamical friction
One term of Chandrasekhar’s formula whose execution frequently varies is the Coulomb logarithm term $\rm ln\left(\Lambda\right)$. $\Lambda$ is related to the ratio of impact parameters $b_{\rm min}/b_{\rm max}$, but writing this quantity in terms of system quantities is difficult (see Just & Peñarrubia (2005) for a discussion of the complexity). In Section 3.2.2 we investigate the role of the $\Lambda$ term, but galpy defaults to:
@@
For the simulations run with dynamical friction, the acceleration is recalculated at every numerical timestep. Each subhalo’s half-mass radius and total mass is used in calculating the acceleration, but like the positions and velocities, these are held constant and not updated throughout the simulation.
Figure 2: Time evolution of average galaxy number density profiles (upper) and differential density profiles (lower) through 10 Gyr of evolution, as shown in the colourbar. The profiles are evolved with DF disabled (on left) and enabled (on right). Haloes are stacked with $M_{200,\mathrm{mean}}$ between $10^{14}$ and $10^{14.5}~{}\mathrm{M}_{\odot}$ (184 haloes), and include all subhaloes.
@@ 2.4 Bootstrapping and fitting
We create galaxy number density profiles for each cluster (see Section 3.2.1 for more discussion). This entails sorting the galaxies by distance from the cluster center and separating them into 32 bins spaced logarithmically from $0.1-10^{0.8}R_{\mathrm{200,mean}}$. The count of galaxies in each bin divided by the volume of that bin is the density profile.
We then take the logarithmic derivative of these profiles, and the minimum of that profile is the splashback radius. To measure the splashback radius in a robust way and quantify uncertainty, we first bootstrap the density profiles and fit them.
For bootstrapping, we provide the density profiles for a given set of clusters (divided by mass) and (sub)set of subhaloes. We bootstrap 1000 times, with 32 sub-bootstraps, and return bootstrapped density profiles and logarithmic density profiles, as well as errors.
To improve quality of the fits, we scale all of the density profiles before bootstrapping. As reference, we use our sample with halo mass between $10^{14}$ and $10^{14.5}$ M and with minimum subhalo stellar mass cut of $10^{7}$ M, and no dynamical friction, since the fitting code works best for this sample. We compute the average density profile for the reference sample and the sample of interest, and find an average scale factor between them. We then multiply all profiles by that factor, while preserving the shape of the density profile. Doing so reduces residual trends where the splashback radius moves inward with increased subhalo mass, and allows for fixed initial free parameters and ranges to be used throughout.
@@
We adjust the fitting boundaries from Appendix B of O’Neil et al. (2021) so that we fit $\rho_{\rm inner}$ at radii $1.20.2<r/R_{200,\rm mean}<$ 1.2, with outer fit extending from $10.02.0<r/R_{200,\rm mean}<$ 10.0. The transition region is fit from $2.01.0<r/R_{200,\rm mean}<$ 2.0, so there is some overlap between the inner and transition regions during fitting. The lack of mergers in our simulation does not have an effect on the splashback radius measurement as the minimum inner radius for fitting is $0.2R_{\mathrm{200,mean}}$.
We discard all fits that do not have a minimum value less than -3 in the differential density profile, as they generally do not trace the splashback curve, as seen by eye. Because our splashback radius values are higher than the typical values measured, (see Figure 5, this is simply because of the way $R_{\mathrm{sp}}$ is defined), the fitting code may preferentially seek low values of the splashback radius. With these adjustments, we are confident in the quality of fits. Visual confirmation suggests that the fits across halo and subhalo mass cuts are accurate.
@@ 3 Results
Figure 1 illustrates the wide variation in how the orbits of galaxies are affected by dynamical friction. Certain galaxies are greatly affected, spiraling into the dense cluster center, while others experience only moderate changes to their orbits. Many seemingly do not undergo any notable changes, such as the galaxies shown more faintly in the background. The galaxies shown in this figure are not representative of the population as a whole, and were selected to demonstrate this variety. Galaxies with properties leading to higher drag forces (galaxies that approach the cluster center closely, are more massive, have higher values of $\Lambda$, etc.) are shown to be more affected by DF. In Figure 1, given the labeled stellar masses, the highlighted galaxies have masses approaching or exceeding the widely used $M_{\mathrm{subhalo}}/M_{\mathrm{halo}}\gtrsim 1\%$ metric, where DF is expected to become influential (Adhikari et al., 2016; More et al., 2016).
@@
These examples illustrate the qualitative impact of dynamical friction on individual galaxy orbits, but to understand how dynamical friction affects our inference of galaxy cluster properties, like the splashback radius, we must expand our analysis to a large sample. The splashback radius is marked by the minimum in the dip of the differential density profile, shown in the bottom panels of Figure 2, which shows the time evolution of galaxy number density and logarithmic density profiles with and without dynamical friction, summed over many clusters. Measurements of $R_{\rm sp}$ are defined relative to $R_{200,\mathrm{mean}}$, because the outskirts of clusters are most self-similar when stacked by $R_{200,\mathrm{mean}}$ (Diemer & Kravtsov, 2014; Lau et al., 2015; Shi, 2016; Umetsu & Diemer, 2017).
@@
The density profile does not evolve dramatically over 10 Gyr, but the differential density profile does exhibit more significant changes with time evolution. The gradual outward creep of the splashback radius, regardless of DF implementation–growing by approximately a factor of two throughout the simulation–does not represent a physical change in the cluster. Because we do not model continuous accretion of mass in our potential, the value of $R_{200,\mathrm{mean}}$ used here is fixed at its initial value from $z=0$. Consequently, as we evolve further into the future, $R_{200,\mathrm{mean}}$ becomes increasingly underestimated relative to the simulated satellite accretion. Further, as shown in O’Neil et al. (2023) and O’Neil et al. (2024), the dip in the differential density profile is created by galaxies with infall times typically over 5 Gyr ago, leading to some small offsets in our simulated properties. As we are generally only concerned with the influence of dynamical friction, we are still able to compare the two simulations to each other and so this shift is not problematic.
@@
With DF enabled, there is a visible overdensity at the center ($\leq 0.1R_{200,\mathrm{mean}}$), that grows with time, which is consistent with the orbital decay that dynamical friction produces. In a real cluster, these galaxies may merge with the central cluster galaxy (e.g. Ostriker & Hausman, 1977); however, this simulation does not have a mechanism for mergers, and therefore these galaxies continue to orbit close to the center. Even though this central impact is clear, the influence of dynamical friction on the splashback region in particular ($\sim R_{\mathrm{200,mean}}$) is not apparent by eye.
@@
To quantify the splashback radius, we bootstrap the density profiles for stacked clusters, as described in Section 2.4. We look at the cluster after 5 Gyr of evolution as at least one crossing time has passed, which we require for the splashback radius to ‘update’ (Figs. 3 through 6 are all at $t=5$ Gyr). Throughout, we divide the clusters into 0.5-dex mass bins for $\log M_{200,\mathrm{mean}}/\mathrm{M_{\odot}}$ in the ranges 13.0 - 13.5, 13.5 - 14.0, 14.0 - 14.5, and above 14.5. Figure 3 shows the bootstrapped differential density profiles, with increasing stellar mass cuts for the subhaloes.
The strength of dynamical friction, as a result of many gravitational interactions, is determined by the subhalo’s total mass rather than its stellar mass. However, throughout our work we present results as a function of stellar mass rather than total mass. We do so because in observation, the limiting factor about whether a galaxy is detected is more closely related to its stellar mass than its total mass. Therefore dividing our galaxy sample by stellar mass allows our simulation to be more easily connected to observational constraints.
The profiles are shown before fitting, producing increased noise – with increased scatter visible as the subhalo sample size shrinks. Even so, Figure 3 visually demonstrates the impact that dynamical friction has on the differential density profiles. Where the profiles appear more blended, dynamical friction has had a weaker effect on the profile. The dip corresponding to the splashback radius is clearly visible, and matches well with the vertical dashed line that identifies the median splashback radius for the sample (post-fitting), such that our visual interpretations here relate well to the quantitative discussion later on.
$M_{*}/\mathrm{M_{\odot}}$ $10^{7}$ $10^{8}$ $10^{9}$ $10^{10}$
$\bar{M}_{\mathrm{subhalo}}/\mathrm{M_{\odot}}$ 1.5 ×$\times 10^{10}$ 5.6 ×$\times 10^{10}$ 1.6 ×$\times 10^{11}$ 4.8 ×$\times 10^{11}$
@@
$\bar{M}_{*}/\mathrm{M_{\odot}}$ 0 1.7 ×$\times 10^{8}$ 2.1 ×$\times 10^{10}$ 1.2 ×$\times 10^{11}$
@@
At less massive stellar mass cuts ($M_{*}>10^{7}~{}\mathrm{M}_{\odot}$) the profiles with and without DF are highly blended. Greater subhalo minimum stellar mass cuts produce increased divergence between the simulations run with and without dynamical friction. We only include minimum mass cuts, rather than i.e. both maximum and minimum masses. Still, because the subhalo mass function can be modeled by a power law with negative slope, most galaxies above a minimum mass cut are close to that minimum mass (e.g. we expect the profile with $M_{*}>10^{7}~{}\mathrm{M}_{\odot}$ to be dominated by galaxies with $M_{*}\sim 10^{7}~{}\mathrm{M}_{\odot}$). As the stellar mass cut increases, both minima move inward (which will be discussed further later), but the profiles with dynamical friction are more significantly shifted towards smaller radii. The changes in the profile, across minimum stellar mass cuts, correspond with the widening separation in the median splashback radius. Notably, this plot examines only the lowest-mass clusters, for which we see the greatest influence of dynamical friction.
After bootstrapping and fitting, we extract values of the splashback radius from the minima of each fitted logarithmic density profile, producing posterior distributions such as the one shown in Figure 4, which shows a histogram of splashback radius distributions for the lowest mass clusters from the bootstrapped and fitted profiles, following the procedure outlined in Section 2.4. That the recorded splashback radii trace a normal distribution indicates that our fitting procedure is stable (e.g. there are no significant outliers and the distribution is not bivariate). Dynamical friction reduces the splashback radius, significantly among high mass galaxies but barely if all galaxies are included (the black curves). As expected, higher galaxy mass cuts correspond with stronger reductions in the splashback radius because of dynamical friction.
However, there is a significant residual trend towards smaller splashback radius as galaxy mass increases, even without dynamical friction. The source of this trend is unclear, but it impacts primarily the lowest-mass clusters (see Figure 5). Because our fit quality for the bootstrapped profiles in this mass bin is high and has been verified by eye, we do not believe it results from a fitting issue. It is possible that the high-mass galaxies trace a different dynamical path than other galaxies, which appears more prominently among the lower-mass clusters.
Our galaxy samples are also greatly affected by the output of the TNG clusters at $z=0$, as they form the initial conditions of our simulations. O’Neil et al. (2022) worked with the same sample of haloes and subhaloes, and found similar, but larger, trends at $z=0$ (0 Gyr). We have examined the “contamination” of our sample at 5 Gyr (i.e. what fraction of galaxies within $2R_{\mathrm{200,mean}}$ at 5 Gyr were inside that radius at the start of our simulation), to see how these galaxies, which primarily trace the splashback radius, might be more or less affected by the starting conditions. However, this contamination fraction is relatively constant across cluster masses. Consequently, we cannot conclude that the $R_{\mathrm{sp}}$ measurements of lower mass clusters are particularly impacted by the simulation starting conditions.
@@
O’Neil et al. (2022) found an $>30\%$ variation in $R_{\mathrm{sp}}$ between stellar mass cuts of $10^{7}$ and $10^{10}~{}\mathrm{M_{\odot}}$ for their lowest mass clusters, and $\sim 20\%$ for the highest mass clusters. We, after 5 Gyr, find a reduced $\sim 15\%$ difference for the lowest mass clusters and little residual discrepancy for the higest mass clusters. Thus, we may simply demonstrate a lingering, but pre-existing, trend, which started more strongly for lower mass clusters.
We are interested in how these trends persist across cluster masses. Figure 5 shows splashback radius distributions, separated by cluster halo mass and subhalo stellar mass. DF reduces the splashback radius, with more prominent effects among subhaloes of higher stellar mass and for clusters of lower mass (<$10^{14}$ M). These results are consistent with the qualitative trends expected of dynamical friction, with high-mass subhaloes in low-mass haloes primarily affected.
In the literature, a variety of approaches, mostly in simulation, have been used to find the minimum mass fraction ($M_{\rm subhalo}/M_{\rm halo}$) that is required for dynamical friction to have a significant impact on the subhalo orbit. We expect that a significant alteration of the orbit will correspond to a change in the splashback radius. Though the subhalo mass is the dynamically relevant quantity, we express our results in terms of stellar mass because it is far more relevant for observations.

### v2

@@ 2 Methods
In this work we use data from the IllustrisTNG (TNG) cosmological simulations to reconstruct simplified realizations of clusters and subhaloes before evolving the subhalo orbits further forwards in time. We use cluster information from TNG because, although we create an idealized simulation to test DF, the initial phase-space distribution of the subhaloes is crucial for replicating the splashback radius. Creating a realistic phase-space distribution of subhaloes organically is very difficult, and it is easy to inadvertently destroy the splashback radius by removing the evolutionary history, since the splashback radius by definition reflects billions of years of the cluster’s evolution (e.g. $R_{\mathrm{sp}}$ is strongly dependent on accretion rate, see Diemer & Kravtsov, 2014; Shin & Diemer, 2023; Towler et al., 2024). These complications are why we use the TNG data instead of creating our own clusters, even if that approach could, in theory, be more flexible. Nevertheless, our methodology for estimating the impact of dynamical friction is relatively insensitive to our initial set-up, as we re-run the same sample twice with dynamical friction “on” and “off”, investigating the differences between the resulting galaxy number density profiles.
The next sections introduce TNG (2.1) and how (sub)haloes in TNG are identified and selected (2.2), before describing the details of our simulation (2.3) and fitting procedure (2.4).
@@ 2.1 IllustrisTNG Simulations
We utilize the $z=0$ snapshot of the largest-volume IllustrisTNG cosmological simulations, as introduced by Nelson et al. (2018); Pillepich et al. (2018a) and publicly released in Nelson et al. (2019). Though the full TNG simulation series includes three box sizes, we use the highest-resolution of the largest volume runs, labeled as TNG300-1. TNG300-1 has a periodic box with side length 302 Mpc, providing a large sample of galaxy clusters. This volume contains $2500^{3}$ dark matter particles with masses of $5.9\times 10^{7}~\mathrm{M_{\odot}}$, and $2500^{3}$ gas particles with target cell masses of $1.1\times 10^{7}~\mathrm{M_{\odot}}$. The dark matter particles have a gravitational softening length of 1.5 kpc in comoving units for $z>1$ and physical units for $z\leq 1$, whereas the gas cells utilize an adaptive softening length (minimum 0.37 kpc).
TNG is built on the magnetohyrodynamical (MHD) moving-mesh code Arepo (Springel, 2010), with an updated version of the Illustris galaxy formation model (Vogelsberger et al., 2013). In the model, gas cells cool through radiative processes and metal line cooling, until they stochastically form stars. As the stars evolve, stellar winds and supernovae return mass, energy, and enriched metals to the gas. Black holes are also included in the simulation; they are seeded in haloes with masses of $1.2\times 10^{6}~\rm M_{\odot}$. TNG updates parts of the Illustris model, including radio mode feedback from AGN (Weinberger et al., 2017), new supernova wind model (Pillepich et al., 2018a), and also includes improvements to the numerical convergence of Arepo (Pakmor et al., 2016).
IllustrisTNG follows a cosmology consistent with the Planck Collaboration et al. (2016a) results. All the runs use cosmological parameters of $\Omega_{m}=\Omega_{dm}+\Omega_{b}=0.3089,\Omega_{b}=0.0486,~\rm and~\Omega_{\Lambda}=0.6911$. Additionally, we use $\sigma_{8}=0.8159,n_{s}=0.9667$. The Hubble constant is set as $H_{0}=100h$ km s-1 Mpc-1, with $h=0.6774$ (Planck Collaboration et al., 2016a). Some runs are dark-matter (gravity) only, and the rest are fully hydrodynamical runs. We utilize the hydrodynamical runs because they contain data about the galaxies around the clusters (e.g. stellar mass).
@@ 2.2 Halo selection and galaxy definition
We follow the halo selection criteria of O’Neil et al. (2021), choosing all haloes of $M_{\mathrm{200,mean}}>10^{13}\mathrm{M_{\odot}}$. Beyond the mass requirement, O’Neil et al. (2022) additionally remove haloes that are within $10R_{\mathrm{200,mean}}$ of a more massive halo, preferentially selecting haloes that are isolated and have not been recently disrupted. This filter is important given that we care about the density profiles outside of the cluster center, and because further simulations of the cluster (see Section 2.3) are based on isolated spherical potentials for each halo. From this process we were left with 1401 haloes at $z=0$. However, we further filter the most massive halo because we identify its density profile as having been significantly disrupted by some less massive haloes within $10R_{\mathrm{200,mean}}$, producing a final sample of 1400 haloes.
@@ 2.3 Simulation details
The subhaloes are integrated in the static gravitational potential of the cluster for 10 Gyr, with the subhalo positions outputted every 1 Gyr for our convenience (all results in the manuscript are shown at t=5 Gyr or t=10 Gyr). The use of a gravitational potential to model the gravitational forces of the cluster is consistent with Adhikari et al. (2016) and is necessary in order to avoid inherently creating dynamical friction. The simulation was run twice from the same starting conditions. In one case, we run with the gravity from a static cluster potential as the only force, in which case there is no dynamical friction implemented. In the second, we run the galpy integration with both the cluster potential and a dynamical friction force, which is described in the next section.
Section 4.1 discusses our motivation for this particular methodology in greater detail. In accordance with our goals for this study, we use a variety of simplifying assumption to effectively isolate dynamical friction, at the expense of a physically realistic cluster.
@@ 2.3.1 Analytical implementation of dynamical friction
One term of Chandrasekhar’s formula whose execution frequently varies is the Coulomb logarithm term $\rm ln\left(\Lambda\right)$. $\Lambda$ is related to the ratio of impact parameters $b_{\rm min}/b_{\rm max}$, but writing this quantity in terms of system quantities is difficult (see Just & Peñarrubia (2005) for a discussion of the complexity). In Appendix B we investigate the role of the $\Lambda$ term, but otherwise use the galpy default of:
@@
For the simulations run with dynamical friction, the acceleration is recalculated at every numerical timestep. Each subhalo’s half-mass radius and total mass is used in calculating the acceleration, but these are held constant and not updated throughout the simulation.
From Equation (1), we expect dynamical friction to reduce the measured splashback radius for high-mass galaxies, particularly in lower-mass clusters, as well as in denser environments, or those with larger $\Lambda$. We investigate primarily the role of cluster and galaxy mass through our main analysis, with a discussion of the Coulomb logarithm in Appendix B.
Figure 2: Time evolution of average galaxy number density profiles (upper) and differential density profiles (lower) through 10 Gyr of evolution, as shown in the colourbar. The profiles are evolved with DF disabled (on left) and enabled (on right). Haloes are stacked with $M_{200,\mathrm{mean}}$ between $10^{14}$ and $10^{14.5}~\mathrm{M}_{\odot}$ (185 haloes), and include all subhaloes.
@@ 2.4 Bootstrapping and fitting
We create galaxy number density profiles for each cluster (see Appendix A for more discussion). This entails binning the galaxies into 32 bins spaced logarithmically from $10^{-1}-10^{0.8}R_{\mathrm{200,mean}}$. The count of galaxies in each bin divided by the volume of that bin produces the density profile.
We then take the logarithmic derivative of these profiles, where the minimum of that profile is the splashback radius. To measure the splashback radius in a robust way and quantify uncertainty, we bootstrap each set of clusters and (sub)set of subhaloes 1000 times, with 32 sub-bootstraps. This process returns bootstrapped density profiles and logarithmic density profiles, as well as errors.
To ensure consistent normalization as we move between cluster masses, we re-scale all of the density profiles before bootstrapping. As reference, we use our sample with halo mass between $10^{14}$ and $10^{14.5}$ M and with minimum subhalo stellar mass cut of $10^{7}$ M, and no dynamical friction, since this sample hosts the largest number of galaxy-mass subhaloes (see Table 1). We compute the average density profile for the reference sample and the sample of interest, and find an average scale factor between them. We then multiply all profiles by that factor, while preserving the shape of the density profile. Doing so reduces spurious trends caused by normalization changes.
@@
We adjust the fitting boundaries from Appendix B of O’Neil et al. (2021) so that we fit $\rho_{\rm inner}$ at radii $0.2<r/R_{200,\rm mean}<1.2$, with outer fit extending from $2.0<r/R_{200,\rm mean}<10.0$. The transition region is fit from $1.2<r/R_{200,\rm mean}<2.4$, so there is some overlap between the transition and outer regions during fitting. The lack of mergers in our simulation thus does not affect the splashback radius measurement as the minimum inner radius for fitting is $0.2R_{\mathrm{200,mean}}$. See Appendix A.2 for further discussion of these regions.
@@ 3 Results
Figure 1 illustrates the wide variation in how the orbits of galaxies are affected by dynamical friction. Certain galaxies are greatly affected, spiraling into the dense cluster center, while others experience only moderate changes to their orbits. Many seemingly do not undergo any notable changes, such as the galaxies shown more faintly in the background. Galaxies with properties leading to higher drag forces (galaxies that approach the cluster center closely, are more massive, have higher values of $\Lambda$, etc.) are shown to be more affected by DF. For example, in Figure 1, given the labeled stellar masses, the highlighted galaxies have masses approaching or exceeding the widely used $M_{\mathrm{subhalo}}/M_{\mathrm{halo}}\gtrsim 1\%$ metric, where DF is expected to become influential (Adhikari et al., 2016; More et al., 2016).
@@
These examples illustrate the qualitative impact of dynamical friction on individual galaxy orbits, but to understand how dynamical friction affects our inference of galaxy cluster properties, like the splashback radius, we must expand our analysis to a large sample. The splashback radius is marked by the minimum in the dip of the differential density profile, as seen in the bottom panels of Figure 2, which shows the time evolution of galaxy number density and logarithmic density profiles with and without dynamical friction, summed over all 185 clusters in the mass bin $10^{14}<M_{\mathrm{200,mean}}/\mathrm{M_{\odot}}<10^{15}$ (see Table 1). Measurements of $R_{\rm sp}$ are defined relative to $R_{200,\mathrm{mean}}$, because the outskirts of clusters are most self-similar when stacked by $R_{200,\mathrm{mean}}$ (Diemer & Kravtsov, 2014; Lau et al., 2015; Shi, 2016; Umetsu & Diemer, 2017).
@@
The density profile does not evolve dramatically over 10 Gyr, but the differential density profile does exhibit more significant changes with time evolution. The gradual outward creep of the splashback radius, regardless of DF implementation–growing by approximately a factor of two throughout the simulation–does not represent a physical change in the cluster. Because we do not model continuous accretion of mass in our potential, the value of $R_{200,\mathrm{mean}}$ used here is fixed at its initial value from $z=0$. As shown in O’Neil et al. (2023) and O’Neil et al. (2024), the dip in the differential density profile is created by galaxies with infall times typically over 5 Gyr ago, leading to some small offsets in our simulated properties. As we are generally only concerned with the influence of dynamical friction, we are still able to compare the two simulations to each other and so this shift is not problematic.
@@
With DF enabled, there is a visible overdensity at the center ($\leq 0.1R_{200,\mathrm{mean}}$), that grows with time, which is consistent with the orbital decay that dynamical friction produces. In a real cluster, these galaxies may merge with the central cluster galaxy (e.g. Ostriker & Hausman, 1977); however, this simulation does not have a mechanism for mergers, and therefore these galaxies continue to orbit close to the center. Still, the influence of dynamical friction on the splashback region ($\sim R_{\mathrm{200,mean}}$) is not immediately apparent.
@@
To quantify the splashback radius, we examine the bootstrapped density profiles for stacked clusters, as described in Section 2.4. We aim to study the cluster after one crossing time has passed, which we require for the splashback radius to ‘update’ without dynamical friction. The cluster crossing time can be approximated as the mass-independent quantity $t_{\mathrm{cross}}=t_{\mathrm{H}}/(5\sqrt{\Omega_{\mathrm{m}}})$ (Diemer et al., 2017), which evaluates to 5.2 Gyr, following the cosmology of O’Neil et al. (2024) Section 3.4. Figs. 3 through 6 thus show our analysis after 5 Gyr of evolution.
Throughout, we divide the clusters into 0.5-dex mass bins for $\log M_{200,\mathrm{mean}}/\mathrm{M_{\odot}}$ in the ranges 13.0 - 13.5, 13.5 - 14.0, 14.0 - 14.5, and above 14.5. We also divide subhalo galaxies by 1 dex minimum stellar mass bins.
The strength of dynamical friction, as a result of many gravitational interactions, is determined by the subhalo’s total mass rather than its stellar mass. However, throughout our work we present results as a function of stellar mass rather than total mass. We do so because in observation, the limiting factor about whether a galaxy is detected is more closely related to its stellar mass than its total mass. Therefore, dividing our galaxy sample by stellar mass allows our simulation to be more easily connected to observational constraints. Additionally, stellar mass is not affected by tidal stripping to the same degree as halo mass (Smith et al., 2016), and we do not have a reasonable stellar evolution model for these galaxies (as we do not track star formation, etc.), making the stellar mass the best proxy for observations. In particular, we follow the stellar mass minimum cuts used in O’Neil et al. (2022) in order to be consistent with their analysis, which was based on IllustrisTNG data. Like O’Neil et al. (2022), we include only minimum mass cuts, rather than both maximum and minimum masses, as an analogue to survey magnitude limits.
Although the dynamical underpinnings of the splashback radius refer to the first particle apocenter, the use of the logarithmic density profile by definition always includes particles which are not tracing out their first apocenter. Nevertheless, the position of the splashback radii from the two methods of measurements correspond (setting aside the asphericality of clusters, which is addressed via stacking; for further discussion see e.g. Diemer & Kravtsov, 2014; Adhikari et al., 2014; Mansfield et al., 2017).
Figure 3 shows the bootstrapped differential density profiles for the lowest cluster mass bin, with increasing stellar mass cuts for the subhaloes. The profiles are shown before fitting, producing increased noise – with increased scatter visible as the subhalo sample size shrinks. Even so, Figure 3 visually demonstrates the impact that dynamical friction has on the differential density profiles. Where the profiles appear more blended, dynamical friction has had a weaker effect on the profile. The dip corresponding to the splashback radius is clearly visible, and matches well with the vertical dashed line that identifies the median splashback radius for the sample (post-fitting), such that our visual interpretations here relate well to the quantitative discussion later on.
$M_{*}/\mathrm{M_{\odot}}$ $10^{7}$ $10^{8}$ $^{9} 1010$$^{10}$
$\bar{M}_{\mathrm{subhalo}}/\mathrm{M_{\odot}}$ 1.5 $\times 10^{10}$ 5.6 $\times 10^{10}$ 1.6 $\times 10^{11}$ 4.8 $\times 10^{11}$
@@
$\bar{M}_{*}/\mathrm{M_{\odot}}$ 0 1.7 $\times 10^{8}$ 2.1 $\times 10^{10}$ 1.2 $\times 10^{11}$
@@
At less massive stellar mass cuts ($M_{*}>10^{7}~\mathrm{M}_{\odot}$) the profiles with and without DF are highly blended. Greater subhalo minimum stellar mass cuts produce increased divergence between the simulations run with and without dynamical friction. As the stellar mass cut increases, both minima move inward (which will be discussed further later), but the profiles with dynamical friction are more significantly shifted towards smaller radii. Notably, this plot examines only the lowest-mass clusters, for which we see the greatest influence of dynamical friction.
After fitting, we extract values of the splashback radius from the minima of each fitted logarithmic density profile, producing posterior distributions such as the one shown in Figure 4, which shows a histogram of splashback radius distributions for the lowest mass clusters. That the recorded splashback radii trace a normal distribution indicates that our fitting procedure is stable (e.g. there are no significant outliers and the distribution is not bivariate). As expected, dynamical friction reduces the splashback radius, significantly among high mass galaxies, though barely when all galaxies are included (the black curves).
@@
We are interested in how these trends persist across cluster masses. Figure 5 shows splashback radius distributions, separated by cluster halo mass and subhalo stellar mass. DF reduces the splashback radius prominently among subhaloes of higher stellar mass and for clusters of lower mass ($10^{14}$ M). These results are consistent with the qualitative trends expected of dynamical friction, with high-mass subhaloes in low-mass haloes primarily affected.